# Resonant amplification of multi-messenger emission in rotating stellar core collapse

### v1

@@ IV Discussion and Conclusions
While our findings are highly promising for multi-messenger observations of CCSNe, further investigation is clearly warranted. In realistic CCSN cores, deviations from axisymmetry arise rapidly; at the rotation rates explored here, non-axisymmetric instabilities may develop and alter the rotational profile, potentially modifying the conditions for resonance. These effects must be rigorously examined using full 3-dimensional (3D) simulations. Furthermore, future studies should probe whether this resonance is sensitive to other progenitor properties, such as initial mass or metallicity, and if it can offer new constraints on the nuclear EOS.
Bearing in mind these caveats, the pronounced, long-lasting modulation of the GW signal induced by resonance reported in this work could represent a breakthrough for multi-messenger detections of CCSNe. Given the substantial GW amplitude enhancement, this resonant signature should remain detectable at distances of $\mathcal{O}(1\,\mathrm{Mpc})$ or, with CE, even $\mathcal{O}(30\,\mathrm{Mpc})$, rendering it a compelling target for both current and next-generation interferometers. Using the CCSN rate estimate [54], the most optimistic source distance of $30\,\mathrm{Mpc}$, and a conservative appraisal that approximately 2% of all the massive stars rotate at rates of around $1\,\mathrm{Hz}$ [55], our findings indicate that future detectors could routinely observe $\sim 1$ CCSN per year.
@@ Acknowledgements.
Acknowledgements. We acknowledge the support through the grants PID2021-127495NB-I00 and PID2021-125485NB-C21 funded by MCIN/AEI/10.13039/501100011033 and by the European Union, and the Astrophysics and High Energy Physics programme of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1) as well as support from the Prometeo excellence programme grants CIPROM/2022/13 and CIPROM/2022/49 funded by the Generalitat Valenciana. MO was supported by the Ramón y Cajal programme of the Agencia Estatal de Investigación (RYC2018-024938-I).

### v2

@@ IV Discussion and Conclusions
While our findings suggest a new effect relevant for multimessenger observations of CCSNe, further investigation is clearly warranted. In realistic CCSN cores, deviations from axisymmetry arise rapidly; at the rotation rates explored here, non-axisymmetric instabilities may develop and alter the rotational profile, potentially modifying the conditions for resonance. These effects must be rigorously examined using full 3-dimensional (3D) simulations. Moreover, 2D simulations may overestimate GW amplitudes compared to 3D [59], which represents a limitation of the present study. Furthermore, future studies should probe whether this resonance is sensitive to other progenitor properties, such as initial mass or metallicity, and if it can offer new constraints on the nuclear EOS.
Bearing in mind these caveats, the pronounced, long-lasting modulation of the GW signal induced by resonance reported in this work could represent a breakthrough for multimessenger detections of CCSNe. Given the substantial GW amplitude enhancement, this resonant signature should remain detectable out to distances of $\sim 1\,\mathrm{Mpc}$ with current detectors. With CE, the detection horizon extends to $\sim 7.4\,\mathrm{Mpc}$. This makes the signal a compelling target for both current and next-generation interferometers.
To provide an estimate of the expected number of detectable events, we approximate the detection rate as
$\mathcal{R}_{\rm det}\;\simeq\;\mathcal{R}_{\rm CCSN}(D<10\,\mathrm{Mpc})\,V(d_{\rm r})\,f_{\rm rot},$ (7)
where $\mathcal{R}_{\rm CCSN}(D<10\,\mathrm{Mpc})$ is the local volumetric core-collapse supernova rate (i.e., within a distance $D<10\,\mathrm{Mpc}$), $V(d_{\rm r})$ is the accessible volume set by the detector horizon distance $d_{\rm r}$, and $f_{\rm rot}$ is the fraction of massive-star progenitors that meet the (model-dependent) conditions for the resonance (e.g. sufficiently rapid presupernova core rotation; namely $\sim 1\,\mathrm{rad/s}$ [60]; see Appendix B). We take $V(d_{\rm r})=4\pi d_{\rm r}^{3}/3$, and adopt $d_{\rm r}\simeq 7.4\,\mathrm{Mpc}$ for CE (as inferred from our SNR analysis). For the local CCSN rate we use the commonly employed empirical estimate of $\sim 1-2$ CCSN per year within $10\,\mathrm{Mpc}$, with uncertainties dominated by incompleteness due to dust extinction and intrinsically faint events (e.g., [61, 62, 63, 64]). That corresponds to $\mathcal{R}_{\rm CCSN}(D<10\,\mathrm{Mpc})\;\approx\;2.4\times 10^{-4}\ {\rm yr}^{-1}\,{\rm Mpc}^{-3}$. Stellar-evolution and population-synthesis studies indicate that only a small fraction of massive-star progenitors are expected to retain presupernova core rotation rates of order $\sim 1\,\mathrm{rad/s}$, with estimates ranging from $f_{\rm rot}\sim 10^{-2}$ up to $\sim 10^{-1}$ in optimistic scenarios involving chemically homogeneous evolution or strong binary interaction (e.g., [65, 40, 60]). As a fiducial, but highly uncertain value one may set $f_{\rm rot}\!=\!0.03$, yielding
$\mathcal{R}_{\rm det}\;\simeq 0012~{\rm yr}^{-1}$ $\left(\frac{f_{\rm rot}}{0.03}\right)\times\left(\frac{d_{\rm hor}}{7.4~{\rm Mpc}}\right)^{3}\times$ (8)
$\left(\frac{\mathcal{R}_{\rm CCSN}(D<10\,\mathrm{Mpc})}{2.4\times 10^{-4}~{\rm yr}^{-1}\,{\rm Mpc}^{-3}}\right),$
i.e., of order one event every $\sim 80$ years.
@@ Acknowledgements.
Acknowledgements. We acknowledge support through the grants PID2021-127495NB-I00 and PID2024-159689NB-C21 funded by MICIU/AEI/10.13039/501100011033 and by FEDER/EU, and the Astrophysics and High Energy Physics programme of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1) as well as support from the Prometeo excellence programme grants CIPROM/2022/13 and CIPROM/2022/49 funded by the Generalitat Valenciana. MO was supported by the Ramón y Cajal programme of the Agencia Estatal de Investigación (RYC2018-024938-I). MC acknowledges the support through the Generalitat Valenciana via the grant CIDEGENT/2019/031. The computations have been performed on servers Lluisvives and Tirant-4 (grant AECT-2025-2-0002) of the Servei d’Informàtica de la Universitat de València and on the Red Española de Supercomputación (RES) on MareNostrum (grants AECT-2025-1-0012 and AECT-2025-2-0006).
Appendix A Resolution study
Two additional high-resolution simulations were performed to assess the dependence of the resonance on spatial resolution. We find that the resonance phenomenon persists, but the associated GW flares become weaker and shorter-lived at higher resolution. This behaviour can be explained by small shifts in the $f$-mode frequency that prevent sustained overlap with the epicyclic frequency, highlighting the strong sensitivity of the resonance to resolution-dependent frequency variations.
All three simulations span the same domain: a radial extent from $0\,\mathrm{cm}$ to $10^{10}\,\mathrm{cm}$, and an azimuthal angle range of $0,\,\pi]$. Model IR uses an innermost grid spacing of $400\,\mathrm{m}$, with $n_{r}=480$ radial zones and $n_{\theta}=128$ azimuthal zones. Model IR-2 refines the innermost spacing to $300\,\mathrm{m}$, with $n_{r}\times n_{\theta}=602\times 256$, while Model IR-3 maintains $300\,\mathrm{m}$ spacing but uses $n_{r}\times n_{\theta}=800\times 512$.
Figure 6: Top panel: GW strains emitted by a source at distance $\mathcal{D}$ for models IR (Black), IR-2 (blue), and IR-3 (orange). Middle panel: evolution of PNS radius (solid lines) and mass (dashed lines). Bottom panel: epicyclic frequency evolution (solid lines) and $f$-mode, computed using [43] (dotted).
Models IR-2 and IR-3 exhibit short flares of enhanced amplitude, lasting a few tens of milliseconds and reaching amplitudes of several tens of centimetres, but both the duration and amplitude are reduced compared to IR (Figure 6 top panel). These flares are more similar in duration and amplitude to models R3 and R4, in which the rotation has been altered (see next section). The bottom panels display the spectrograms. In higher-resolution runs the $f$-mode frequency, indicated by the darker band of increasing frequency, is shifted upward by $\sim 100\,\mathrm{Hz}$, so the fundamental epicyclic frequency (solid blue lines) and its maximum (blue shading) do not remain aligned with the $f$-mode long enough to drive a strong resonance, producing only short, weak flares.
Consistency across resolutions is further confirmed in the bottom two panels of Figure 6. PNS quantities for IR and IR-2 agree within 5%. Model IR-3 shows a slightly larger and more oblate PNS, due to its faster rotation. The epicyclic frequencies agree within $100\,\mathrm{Hz}$ overall, but in the first $0.3\,\mathrm{s}$ post-bounce, when the initial resonant bursts appear in IR, the higher-resolution models yield slightly higher values. This may delay the crossing with the $f$-mode beyond the computed times or happen in shorter periods, thereby damping the resonance.
As in many other resonant phenomena in physics (e.g., a driven, slightly damped pendulum or the propagation of seismic waves with anelastic attenuation), damping implies that a higher quality factor $Q$ results in a narrower frequency range over which true resonance occurs. The resolution study presented here suggests that the resonance observed has a relatively high $Q$, so even small detuning between the epicyclic frequency and the PNS $f$-mode frequency significantly reduces the resonant amplitude. Consequently, time-dependent factors that slightly alter the conditions governing either the $f$-mode or the epicyclic frequency can induce partial quenching or intermittency in the resonant response.
This argument explains how changes in resolution—leading to the mild dynamical differences discussed in previous paragraphs—can cause a detuning between the $^{2}f$ and $f_{\rm epi}$ modes and, accordingly, a significant reduction in the GW amplitude. This result should not be interpreted as a lack of physical relevance of the identified resonance. Rather, it suggests that very slight modifications to the initial conditions—specifically, to the central nuclear rotational rate $\Omega_{\rm c}$—can trigger large-amplitude GWs in these models (as observed for $\Omega_{\rm c}=1\,\mathrm{rad/s}$). The substantial computational cost of higher-resolution models prevents a more detailed exploration to confirm this point. However, this interpretation is further supported by additional simulations based on different presupernova stellar models (not shown in this work) that exhibit similar resonant behaviour.
We also attribute to the high $Q$ the partial quenching of the resonance observed between the two main flares in model IR. The low-activity period coincides with a phase during which extended equatorial inflows strike the PNS surface. Most of the high–specific-angular-momentum material impacting the PNS is not assimilated (as indicated by the absence of a clear increase in PNS mass); instead, it rebounds and is eventually expelled through the polar regions into the collimated supernova ejecta. We tentatively identify the random variations induced by these inflows on the PNS as a plausible source of detuning between the $f$-mode and epicyclic frequencies.
In summary, the resonance is robust to changes in resolution, but its GW signature depends sensitively on the relative values of the $f$-mode and epicyclic frequencies. Further high-resolution simulations are required to quantify the enhancement produced by this resonance and to constrain the parameter space in which it arises.
Appendix B Rotation dependence
Two additional simulations exploring the rotation parameter space were carried out to test how the resonance depends on the initial rotation rate. These correspond to model R3, initialized with three times the rotation rate of the progenitor, and model R4, with a fourfold increase. These models are configured with nuclear rotational rates near the value at which we observed resonant behaviour, specifically $\Omega_{\rm c}=0.87$, 1, and $1.16\,\mathrm{rad/s}$. The results show that the resonance is highly sensitive to rotation: increasing or decreasing the initial rotation changes both the timing and the character of the GW signal.
Figure 7: Same as Figure 6 but for models IR, R3, and R4, in which the $f$-mode has been replaced with $T/|W|$.
The top panel of Figure 7 compares the GW strains of models R3, R4, and IR. The reference model IR exhibits a continuous, sustained resonance over $0.4{-}0.8\,\mathrm{s}$. Model R3, instead, shows a strong but short-lived enhancement lasting about $100\,\mathrm{ms}$, beginning at $0.35\,\mathrm{s}$, followed by a quiescent phase until $0.9\,\mathrm{s}$. Model R4 produces yet another pattern: a sequence of intermittent flares of elevated GW amplitude, each lasting only a few tens of milliseconds. Together, these comparisons demonstrate that the resonance peaks for a specific rotation, while weakens or fragments for increased or decreased spin.
The middle panel of Figure 7 shows the PNS mass (dashed lines) and radius (solid lines). The three models produce very similar PNS structures, with only a small mass offset: model R3 is more massive by about $\sim 0.05\,\mathrm{M_{\odot}}$, while the radii remain nearly identical.
The bottom panel of Figure 7 compares the epicyclic frequency (solid lines) and the ratio of rotational to gravitational energy, $T/|W|$ (dotted lines). In the slower-rotating models (IR and R3), the epicyclic frequencies are nearly identical between $0.2{-}0.6\,\mathrm{s}$. However, in model R3, $T/|W|$ gradually increases and eventually surpasses that of IR, even though IR began with faster rotation. After resonance is triggered, the value of $T/|W|$ in IR stabilizes, while it continues to rise in R3. In the fastest-rotating case, R4, both the epicyclic frequency and $T/|W|$ are consistently higher by about $\sim 100\,\mathrm{Hz}$ and $0.005$, respectively. These differences indicate that changes in the PNS shape and rotation rate can shift the frequency of the $f$-mode, altering the conditions required for the epicyclic frequency to maintain resonance. As a result, sustained enhancements of the GW signal become harder to achieve at higher rotation rates.
The sensitivity to the rotational rate must be interpreted with care. Changes in numerical resolution propitiate a slight detuning between the $f$-mode and epicyclic frequencies, driven by mild differences in the dynamics. We suspect that resolution-dependent effects may lead to conditions in which the resonance is amplified for different initial values of $\Omega_{\rm c}$. The substantial computational resources required to verify this hypothesis compel us to defer a detailed study to future work.
In summary, the resonance is found within a relatively narrow range around $\Omega_{\rm c}\approx 1\,\mathrm{rad/s}$ under the idealized conditions of our models, where rotation is imposed on an initially non-rotating stellar structure. Combined with the uncertainties in the actual rotational profiles predicted by stellar evolution models, this makes it very difficult to estimate the fraction of massive, rotating stars in which conditions are favourable for the development of the resonance. Such conditions may occur in, say, $\sim 10\%$ of the tail of faster rotators (i.e., with equatorial projected rotational speeds $v\sin i>150\,\rm km\,s^{-1}$; [60]). This implies that only about $1\%$ of all massive stars could host favourable conditions for the resonance discussed in this work.